Without expanding completely, find the sixth term in the expansion of the expression.
((1/c)+(c^2/2))^8

Answer by stanbon(75887)

find the sixth term in the expansion of the expression.
((1/c)+(c^2/2))^8
----
8C3(1/c)^3*(c^2/2)^5
-----
= 56(1/c^3)(c^10/32))
----
= (56/32)c^7
----
= (7/4)c^7
================
Cheers,
Stan H.